OpenStudy (anonymous):

First estimate the answer. Then subtract to find the exact answer. 28 7/9 - 5 5/6 = 28 7/9 rounds to ? 5 5/6 roundd to ?

OpenStudy (wolfe8):

Can you use a calculator? You should be able to. When you have a fraction, it means the numerator divided by the denominator. A mixed number like the ones you have, with an integer in front of a fraction, means the integer itself. So 28 7/9 is 28 + 7/9

OpenStudy (anonymous):

but i have to round them first and then solve and then dont round them and solve again

OpenStudy (wolfe8):

I'm sorry I don't really understand the question. Estimate first? If you want them rounded off, like I said, divide the numerator by the denominator.

OpenStudy (anonymous):

yeah i have to round it off but i already did that part now i just need to solve 28 7/9 - 5 5/6 = ?

OpenStudy (wolfe8):

Ah so when rounded off that is the estimate? So if you want the exact value it has to remain in fraction form, because that is why we even have fractions; to represent divisible numbers with a lot of or endless decimal places. So subtract them when in fractions first. After that, I'm not sure if you are required to get the decimal form of the answer.

OpenStudy (anonymous):

it say type a integer , proper fraction, or mixed number and it says to simplify it

OpenStudy (wolfe8):

Then it has to be in fraction form in the end :)

OpenStudy (anonymous):

yes but i dont know how to solve it

OpenStudy (wolfe8):

Well when written correctly, it is\[28\frac{ 2 }{ 9 } -5\frac{ 5 }{ 6 }\] You can just subtract the whole numbers; 28-5 Then you will have to find a common denominator for 2/9 and 5/6. Hint: it should be a common multiple of 6 and 9 ;) Then you multiple each fraction with whatever values were needed to obtain that common denominator.(They should be different for each fraction) And then you can subtract as usual. This might help http://www.mathsisfun.com/fractions_subtraction.html

OpenStudy (anonymous):

do you know how to do this?

OpenStudy (wolfe8):

Yes but can you not do it? If the link doesn't help I will show you myself.

OpenStudy (anonymous):

i understand it but i got it wrong from the link i did the steps from the link and got it wrong.

OpenStudy (wolfe8):

Alright so you have \[(28-5)(\frac{ 2 }{ 9 }-\frac{ 5 }{ 6 })\]\[25\left( \frac{ 2\times2 }{9\times2 } -\frac{ 5\times3 }{6\times3 } \right)\]\[25\left( \frac{ 4-15 }{ 18 } \right)\]

OpenStudy (anonymous):

so the answer would be 25 11/18

OpenStudy (wolfe8):

Ah but notice that 4-15 will give you a negative number.

OpenStudy (anonymous):

yeah i put it and got it wrong

OpenStudy (wolfe8):

It should be -25 11/18

OpenStudy (anonymous):

it said its wrong

jimthompson5910 (jim_thompson5910):

28 7/9 rounds to 29 5 5/6 rounds to 6 so 28 7/9 - 5 5/6 is estimated by 29 - 6 = 23

OpenStudy (wolfe8):

...what the.. 28-5 = 23...

jimthompson5910 (jim_thompson5910):

\[\large 28 \frac{7}{9} - 5 \frac{5}{6}\] \[\large 28 \frac{7*2}{9*2} - 5 \frac{5*3}{6*3}\] \[\large 28 \frac{14}{18} - 5 \frac{15}{18}\] \[\large \frac{28*18+14}{18} - \frac{5*18+15}{18}\] \[\large \frac{504+14}{18} - \frac{90+15}{18}\] \[\large \frac{518}{18} - \frac{105}{18}\] \[\large \frac{518-105}{18}\] \[\large \frac{413}{18}\] \[\large \frac{396+17}{18}\] \[\large \frac{396}{18}+\frac{17}{18}\] \[\large 22+\frac{17}{18}\] \[\large 22\frac{17}{18}\] So that means \[\large 28 \frac{7}{9} - 5 \frac{5}{6} = 22\frac{17}{18}\]

jimthompson5910 (jim_thompson5910):

so 23 is a pretty good estimate if you ask me

OpenStudy (anonymous):

thanks so much @jim_thompson5910 can you help me with like 3 more please.

OpenStudy (anonymous):

these next 3 i got wrong

jimthompson5910 (jim_thompson5910):

sure post them as new questions